# CoRanking: Collaborative Ranking with Small and Large Ranking Agents

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated superior listwise ranking performance. However, their superior performance often relies on large-scale parameters (e.g., GPT-4) and a repetitive sliding window process, which introduces significant efficiency challenges. In this paper, we propose CoRanking, a novel collaborative ranking framework that combines small and large ranking models for efficient and effective ranking. CoRanking first employs a small-size reranker to pre-rank all the candidate passages, bringing relevant ones to the top part of the list (e.g., top-20). Then, the LLM listwise reranker is applied to only rerank these top-ranked passages instead of the whole list, substantially enhancing overall ranking efficiency. Although more efficient, previous studies have revealed that the LLM listwise reranker have significant positional biases on the order of input passages. Directly feed the top-ranked passages from small reranker may result in the sub-optimal performance of LLM listwise reranker. To alleviate this problem, we introduce a passage order adjuster trained via reinforcement learning, which reorders the top passages from the small reranker to align with the LLM’s preferences of passage order. Extensive experiments on three IR benchmarks demonstrate that CoRanking significantly improves efficiency (reducing ranking latency by about 70%) while achieving even better effectiveness compared to using only the LLM listwise reranker.
@@
CoRanking: Collaborative Ranking with Small and Large Ranking Agents
@@
Figure 1: (A) Apply LLM listwise reranker to rerank 100 passages based on sliding window strategy. (B) Utilize a small reranker to pre-rank 100 passages and then apply LLM reranker to rerank top-20 passages. (C) After the reranking process of small reranker, we first use a passage order adjuster to adjust the top-20 passages and then apply the LLM reranker to rerank top-20 passages.
@@ 1 Introduction
In recent years, many studies Liu et al. (2024c); Sun et al. (2023); Zhu et al. (2023) have demonstrated strong zero-shot ranking capabilities of large language models (LLMs). A typical approach is listwise ranking which feeds the query and a list of passages into LLM and instructs it to output a reranked list of passage IDs in descending order of relevance. This approach has been widely adopted by many studies, outperforming many ranking approaches Qin et al. (2023); Liang et al. (2022); Sachan et al. (2022) and establishing state-of-the-art performance on major IR benchmarks including TREC Craswell et al. (2020b) and BEIR Thakur et al. (2021) through implementations like RankGPT Sun et al. (2023) using GPT-4.
@@
Despite its effectiveness, listwise ranking faces critical efficiency challenges (Figure 1A), primarily stemming from two factors. First, its state-of-the-art performance relies on massive LLMs (e.g., Qwen2.5-72B Yang et al. (2024) and GPT-4), which introduces significant computational overhead and results in the efficiency issue. Second, to achieve a trade-off between effectiveness and efficiency, listwise ranking usually utilize the sliding window strategy. With fixed window size and step size, this strategy bubbles the relevant passages that are initially lowly ranked to the top. However, the interdependence between windows hinders the parallelization of sliding windows, which results in many repetitive and sequential ranking of sliding window (e.g., totally 9 sliding windows if window size and step size are set as 20 and 10 when reranking 100 passages) and creates an efficiency bottleneck (see Figure 1 (A)).
@@
The sliding window mechanism of LLM reranker primarily serves to elevate initially low-ranked relevant passages into the final reranking window (e.g., top-20). We believe that this repetitive and inefficient process could be replaced by a smaller reranker: By utilizing a smaller reranker to pre-rank the list of passages, the listwise LLM reranker only needs to rerank the passages in the final window. This hybrid approach could dramatically reduce computational costs and enhance the ranking efficiency.
@@
However, existing small rerankers are primarily fine-tuned from relevance labels (human labels Nogueira and Cho (2019) or teacher rankings Pradeep et al. (2023b); Liu et al. (2024c)). We argue that directly feeding the top-ranked passages of these reranker into an LLM listwise reranker may lead to sub-optimal performance (see Figure 1 (B)). This is because the order of input passages significantly impacts the effectiveness of the listwise LLM reranker Sun et al. (2023); Tang et al. (2024), which stems from LLMs’ pre-training biases toward the positions of input tokens. It indicates that the performance of the LLM listwise reranker has a complex relationship with the order of the input passages, rather than solely depending on the relevance ranking of the passages. Thus, the passage order given by small reranker can hardly aligns with the order preferences of LLMs. To address this issue, we introduce a passage order adjuster to rerank the top-ranked passages from the small reranker so that they align with the order preference of the LLM listwise reranker (see Figure 1 (C)).
@@
In this paper, we propose a collaborative ranking framework CoRanking that integrates small and large ranking agents to perform passage ranking in a multi-stage manner. In the first stage, we train a small listwise reranker (SLR) to pre-rank all the passages to elevate the relevant passages to top part of list (e.g., top-20). Different from existing methods Pradeep et al. (2023a, b) that solely use teacher-generated ranking lists as training labels, we propose a human-label-enhanced method to improve the ranking quality. In the second stage, we train a passage order adjuster (POA) using the reinforcement learning algorithm DPO. The POA reorders the top-ranked passages to align with LLM’s order preference. In addition, we design a significance-aware selection strategy $\text{S}^{3}$ to build high-quality preference pairs for DPO training. In the third stage, an LLM listwise reranker is used to further rerank the top-ranked passages, enhancing the overall ranking effectiveness. We conduct extensive experiments on diverse IR benchmarks and the results indicate that our CoRanking framework achieves 70% faster ranking time and even better ranking effectiveness than pure LLM listwise reranker.
@@
(1) We propose a ranking framework CoRanking that enables small and large ranking agents to collaborate with each other, thereby achieving efficient and effective passage ranking.
(2) We introduce a novel passage order adjuster (POA) that aligns the the order of reranked passages from small reranker with the preference of large listwise reranker.
(3) We design the $\text{S}^{3}$ strategy which build valuable training samples for the DPO algorithm during POA’s training, improving the effectiveness of preference alignment.
Figure 2: Our collaborative ranking framework CoRanking consists of three stages. In the first stage, SLR reranks top-100 passages based on sliding window strategy. In the second stage, POA reranks the top-20 passages to align with the positional preference of LLR. In the final stage, LLR is utilized to rerank the top-20 passage to obtain the final ranking list. The training details of SLR and POA are shown in yellow and cyan parts respectively.
@@ 2 Methodology
In this section, we present a collaborative ranking framework CoRanking that integrates both small and large passage rerankers. As illustrated in Figure 2, the framework comprises three key components: (1) a small listwise reranker (SLR) for preliminary passage reranking, (2) a passage order adjuster (POA) that aligns the order of SLR’s top-ranked passages with the LLM listwise reranker’s preference, and (3) an LLM listwise reranker (LLR) for final ranking. We begin by formalizing the passage ranking task, then elaborate on each component’s design.
@@ 2.1 Task Definition
The objective of passage ranking is to rerank a list of retrieved passages $[p_{1},\ldots,p_{N}]$ based on their relevance to a query $q$. The listwise ranking approach takes both $q$ and a set of passages as input, and generates a reranked sequence of passage IDs (e.g., [4][2]…). Due to LLMs’ input length constraints, existing listwise ranking approaches typically employ a sliding window strategy that iteratively processes passage subsets. Specifically, this strategy applies a window of size $w$ that moves from the end to the beginning of the passage list with a step size $s$, dynamically promoting relevant passages to higher positions. Following previous studies Sun et al. (2023); Liu et al. (2024c), we set the total number of passages $N$ to 100, with window size $w$ and step size $s$ set to 20 and 10, respectively.
2.2 Stage1: Small Listwise Reranker (SLR)
@@
To mitigate the latency and computational overhead of the LLM listwise reranker (LLR), we propose using a small listwise reranker (SLR) to pre-rank the entire passage list. The SLR aims to elevate relevant passages into the final sliding window (i.e., the top-20 passages), thereby reducing the number of LLR’s sliding window iterations. Here we adopt listwise training because we find that it performs better than pointwise training and the SFT-trained SLR naturally serves as the initialization for our listwise passage order adjuster, which will be detailed in Section 2.3.
@@
Training Label. Conventional listwise reranker training typically employs knowledge distillation Pradeep et al. (2023a, b), which uses a strong teacher model (e.g., GPT-4) to rerank BM25-retrieved top-20 passages, with the teacher-generated rankings serving as supervised fine-tuning (SFT) labels. However, the teacher model can not perform consistently well across all queries (e.g., long-tail queries), which can result in suboptimal performance on these queries. To further improve the SFT label, we propose a human-label-enhanced ranking construction (HRC) method that improves teacher-generated rankings with human-label.
@@
Specifically, we first iteratively rerank BM25-retrieved top-20 passages $M$ times using the teacher model in a listwise manner, where the $i$-th reranking input is the output ranking from the $(i-1)$-th iteration. This progressive reranking approach, motivated by prior studies Pradeep et al. (2023b, a), enables gradual improvement of ranking lists. We select the iteration where human-annotated relevant passages are ranked at the top (NDCG=1) as the gold ranking.
@@
Training Loss. After obtaining the gold ranking, we concatenate the query $q$ and the candidate passages $[p_{1},\ldots,p_{N}]$ together as the training input $x$ and optimize our SLR by minimizing the standard language modeling loss $\mathcal{L}$:
@@
$x$ $=q\circ\text{[1]}\circ p_{1}\circ\text{[2]}\circ p_{2}\ldots\text{[N]}\circ p_{N}$ (2)
@@
2.3 Stage2: Passage Order Adjuster (POA)
While our SLR is trained to rank relevant passages at the top of the list, the order of top-ranked passages (i.e., the top-20 passages) may not align well with the positional preferences of LLR. To address this mismatch, we employ reinforcement learning (RL) algorithm to train a passage order adjuster POA based on LLR’s feedback to rerank these passages. We adopt the Direct Preference Optimization (DPO) Rafailov et al. (2023) algorithm for its optimization stability. The following sections elaborate on the construction of DPO training data and model optimization.
Training Data. The training data of DPO consists of a series of prediction pairs. Each prediction pair contains a winner and a loser prediction, namely passage ID lists. The reward $R$ for each prediction $L$ is computed by feeding the corresponding passage list to LLR and evaluating the ranking metric $M$ of LLR’s output ranking $L^{\prime}$ based on human-annotated label $y$:
@@ 2.3 Stage2: Passage Order Adjuster (POA)
We apply our trained SLR to rerank BM25-retrieved top-100 passages for training queries and use the top-20 reranked passages as training candidate passages, which aims to align with the input passage distribution encountered by POA during inference. To generate the preference pairs for training POA, we first sample $m$ predictions from SLR by taking the candidates as the input, then evaluate these predictions through Equation 3. After that, the highest-reward list is selected as the winner prediction, while others serve as loser predictions. Next, we pair the winner prediction with all the loser predictions to construct initial training preference pairs.
@@
Optimization. Given constructed preference pairs, we optimize POA using the following DPO objective:
@@
where $x$ represents the training input whose format is the same as Equation 1 and $y_{w}$, $y_{l}$ represent the winner prediction and loser prediction, respectively. $\pi_{\theta}$ is the policy model to be optimized and $\pi_{f}$ is the original policy model which is initialized with our trained SLR. The frozen $\pi_{f}$ acts as a regularizer to maintain generation stability during optimization and $\beta$ balances preference alignment and regularization towards $\pi_{f}$.
2.4 Stage3: LLM Listwise Reranker (LLR)
After the Passage Order Adjuster (POA) adjusts the top-20 passages, we use the LLM Listwise Reranker (LLR) to rerank these 20 passages. This avoids LLR’s inefficient sliding window operations on all 100 passages, significantly improving ranking efficiency. The specific prompt template for LLR is provided in Appendix A, which is consistent with previous studies Pradeep et al. (2023b); Liu et al. (2024c).
@@ 3.1 Setting
We evaluate our framework on three established information retrieval benchmarks: TREC DL Craswell et al. (2020b), BEIR Thakur et al. (2021), and BRIGHT Su et al. (2024). For the TREC DL benchmark, we adopt the TREC Deep Learning 2019 (DL19) Craswell et al. (2020b) and 2020 (DL20) Craswell et al. (2020a) datasets derived from MS MARCO V1. The BEIR benchmark Thakur et al. (2021) is selected to evaluate cross-domain generalization capabilities. We choose three representative datasets: TREC-Covid, Robust04, and Trec-News. BRIGHT Su et al. (2024) is a reasoning-intensive IR benchmark whose relevance assessment transcends simplistic keyword matching or semantic similarity, instead demanding intentional and deliberate reasoning. We choose Economics, Earth Science and Robotics datasets for evaluation. The detailed descriptions of each dataset will be provided in Appendix B.
@@
Table 1: Results (NDCG@10) on TREC, BEIR and BRIGHT benchmarks. The “Pointwise”, “Sliding”, ’Full’ and “Collaborative” in “Strategy” column represent pointwise ranking, sliding window strategy, full ranking Liu et al. (2024c) and collaborative ranking, respectively. The best and second-best results for “Single Reranker” part and “Collaborative Rerankers” part are indicated in bold and underlined, respectively. “Avg.” represents the averaged result of all 8 datasets.
Implementation Details
As for the supervised fine-tuning of our SLR, we use Qwen2.5-3B-Instruct as the backbone model. The training queries consists of totally 2k queries sampled from MS MARCO training set. We utilize Qwen2.5-72B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-GPTQ-Int4 as our teacher model to SFT our SLR and also employ it as the LLR model throughout this study unless otherwise specified. Unlike previous studies Pradeep et al. (2023b); Liu et al. (2024c) that adopted GPT-series models (e.g., GPT-4o) as the teacher model, we select Qwen2.5-72B-Instruct as it demonstrates comparable and even better performance to GPT-4o across most datasets while being more cost-effective. Nevertheless, we have included additional experimental results using GPT-4o as the LLR model in Section 3.6. We train SLR for 4 epochs with a learning rate of 5e-6 and batch size of 1. As for the RL training of POA, we use our trained SLR for model initialization. We set the threshold $\mu$ as 0.4 and generate about 12k preference pairs based on 2k queries for DPO algorithm. The batch size, learning rate and the hyper-parameter $\beta$ of DPO are set as 1, 1e-6 and 0.4 respectively.
Baselines
As for single reranker, we compare our trained SLR with a series of state-of-the-art supervised rerankers, including monoBERT (340M) Nogueira and Cho (2019), monoT5 (3B) Nogueira et al. (2020), RankT5 (3B) Zhuang et al. (2023b), RankVicuna (7B) Pradeep et al. (2023a), RankZephyr (7B) Pradeep et al. (2023b) and $\textbf{RankMistral}_{\text{100}}$. The first three are pointwise rerankers trained from human labels in MS MARCO dataset. The last three are listwise rerankers which are distilled from GPT-3.5, GPT-4 and GPT-4o, respectively. To ensure more fair comparison with SLR in model architecture and training data, we introduce baseline Qwen-pointwise which is a pointwise reranker based on Qwen2.5-3B-Instruct. This model concatenates query-passage pair as input and outputs a relevance score. We utilize the same gold ranking label as SLR and optimize the model with RankNet loss, which is the same as what  Sun et al. (2023) did. We also report the ranking performance of our SLR and LLR which reranks all the passages using sliding window strategy.
As for collaborative rerankers, we introduce three baselines, including monoT5 (3B) + LLR, RankT5 (3B) + LLR and Qwen-pointwise (3B) + LLR. The three baselines first use corresponding small reranker (3B) to rerank all the passages and then apply LLR to rerank top-20 passages.

### v3

# Wenhan Liu1, Xinyu Ma2, Yutao Zhu1, Lixin Su2, Shuaiqiang Wang2 Dawei Yin2 Zhicheng Dou1

@@ Abstract
Listwise ranking based on Large Language Models (LLMs) has achieved state-of-the-art performance in Information Retrieval (IR). However, their effectiveness often depends on LLMs with massive parameter scales and computationally expensive sliding window processing, leading to substantial efficiency bottlenecks. In this paper, we propose a Collaborative Ranking framework (CoRanking) for LLM-based listwise ranking. Specifically, we strategically combine an efficient small reranker and an effective large reranker for collaborative ranking. The small reranker performs initial passage ranking, effectively filtering the passage set to a condensed top-k list (e.g., top-20 passages), and the large reranker (with stronger ranking capability) then reranks only this condensed subset rather than the full list, significantly improving efficiency. We further address that directly passing the top-ranked passages from the small reranker to the large reranker is suboptimal because of the LLM’s strong positional bias in processing input sequences. To resolve this issue, we propose a passage order adjuster learned by RL that dynamically reorders the top passages returned by the small reranker to better align with the large LLM’s input preferences. Our extensive experiments across three IR benchmarks demonstrate that CoRanking achieves superior efficiency, reducing ranking latency by approximately 70% while simultaneously improving effectiveness, compared to the standalone large reranker.
@@
[Uncaptioned image] CoRanking: Collaborative Ranking with Small and Large Ranking Agents
@@
Figure 1: (A) Large Listwise Reranker: Apply a large listwise reranker to rerank 20 passages from 100 candidates based on a sliding window strategy. (B) Naive Collaboration: Pre-rank 100 passages with a small listwise reranker using a sliding window strategy, then rerank the top 20 with a large listwise reranker. (C) CoRanking: Pre-rank 100 passages with a small reranker using a sliding window strategy, adjust the order of the top 20 via the passage order adjuster, then rerank them with the large listwise reranker. “Large/Small Reranker Calls” represents the number of sliding windows for listwise ranking.
@@ 1 Introduction
In recent years, many studies Liu et al. (2024c); Sun et al. (2023); Zhu et al. (2023) have demonstrated strong zero-shot passage ranking capabilities of large language models (LLMs). A typical approach is listwise ranking, which feeds the query and a list of passages into an LLM and instructs it to output a ranked list of passage IDs in descending order of relevance. This approach has established state-of-the-art performance Sun et al. (2023) on major IR benchmarks, including TREC Craswell et al. (2020b) and BEIR Thakur et al. (2021).
@@
Despite its effectiveness, LLM listwise passage ranking faces critical efficiency challenges, as illustrated in Figure 1(A), primarily stemming from two factors: First, its state-of-the-art performance relies on LLMs with massive parameters (such as Qwen2.5-72B or GPT-4), which introduces significant computational overhead and results in an efficiency issue. Second, due to the limitation of the context length of LLMs, existing listwise ranking methods typically employ a sliding window strategy Sun et al. (2023). With a fixed window size (e.g., 20 passages) and step size (e.g., 10 passages), this strategy bubbles the relevant passages that are initially lowly ranked to the top. However, the interdependence between windows hinders the parallelization of sliding windows, which results in many repetitive and sequential rankings of sliding windows (e.g., totally 9 sliding windows if window size and step size are set as 20 and 10 when ranking the top 20 from 100 passages) and creates an efficiency bottleneck (especially when employing LLMs with massive parameters).
@@
The sliding window strategy of LLM reranker primarily serves to elevate initially low-ranked relevant passages into the final reranking window (e.g., top 20). We believe that this repetitive and inefficient process could be replaced by a small listwise reranker. [^1] In this work, we consider models with parameters less than or equal to 3B as small rerankers. By utilizing a small listwise reranker to pre-rank the whole list of passages (e.g., 100 passages) using a sliding window strategy, the large listwise reranker, which has stronger ranking capability, only needs to rerank the passages in the final window (i.e., containing the top 20 passages). This hybrid approach could dramatically reduce computational costs and improve efficiency.
@@
However, existing small listwise rerankers are mainly trained on golden lists whose passages are ranked solely based on relevance Pradeep et al. (2023b, a). We argue that directly feeding the top-ranked passages returned by these small listwise rerankers into a large listwise reranker may lead to sub-optimal performance (see Figure 1 (B)). This is because the effectiveness of a listwise reranker has a complex relationship with the order of input passages, rather than being determined exclusively by their relevance, as indicated by existing studies Sun et al. (2023); Tang et al. (2024). For example,  Tang et al. (2024) reveal that relevant passages appearing in the middle of the list tend to be ranked at the bottom by the listwise reranker. This stems from LLMs’ pre-training biases toward the positions of input tokens Liu et al. (2024a). While the passage order generated by the small listwise reranker, which is trained using relevance-based ranking, may deviate from the large listwise reranker’s order preference. To address this issue, we introduce a passage order adjuster (also a small listwise reranker) to rerank the top-ranked passages from the small listwise reranker so that they align with the order preference of the large listwise reranker (see Figure 1 (C)).
@@
More specifically, we propose a collaborative ranking framework, namely CoRanking, that integrates small and large ranking models to perform passage ranking in a multi-stage manner. In the first stage, we train a small listwise reranker (SLR) to pre-rank all the passages to elevate the relevant passages to the top part of the list (e.g., top-20). Different from existing methods Pradeep et al. (2023a, b) that solely use teacher-generated ranking lists as training labels, we propose a human-label-enhanced method to improve the ranking quality. In the second stage, we train a passage order adjuster (POA) using the reinforcement learning algorithm DPO. The POA reorders the top-ranked passages to align with the large listwise reranker’s order preference. In addition, we design a significance-aware selection strategy $\text{S}^{3}$ to build high-quality preference pairs for DPO training. In the third stage, a large listwise reranker is used to further rerank the top-ranked passages, enhancing the overall ranking effectiveness. Extensive experiments on diverse IR benchmarks demonstrate that our CoRanking framework achieves a 70% reduction in ranking latency while maintaining superior effectiveness, compared to a standalone large listwise reranker.
@@
(1) We propose a ranking framework, CoRanking, that enables small and large LLMs to collaborate, thereby achieving efficient and effective passage ranking.
(2) We propose a novel passage order adjuster (POA) that aligns the output passages of the small listwise reranker with the order preference of the large listwise reranker. We further design the $\text{S}^{3}$ strategy, which builds valuable preference pairs for POA’s training.
(3) We conduct extensive experiments on diverse IR benchmarks, and the results demonstrate that our CoRanking achieves a significant efficiency improvement compared to a standalone large listwise reranker while exhibiting superior performance.
Figure 2: The “Inference” part shows how our CoRanking works with both large and small listwise rerankers. As for training of the small listwise reranker, we design a human-label-enhanced ranking construction method to generate gold rankings for supervised fine-tuning. As for training the passage order adjuster, we proposed the $\text{S}^{3}$ strategy to obtain high-quality preference pairs for DPO reinforcement learning.
@@ 2 Methodology
In this section, we present a collaborative ranking framework CoRanking that integrates both small and large passage rerankers. As illustrated in Figure 2, the framework comprises three key components: (1) a small listwise reranker, (2) a passage order adjuster, and (3) a large listwise reranker. Next, we will introduce the definition of the ranking task, the inference process of CoRanking, and the training details.
@@ 2.1 Task Definition
The aim of passage ranking is to rerank a list of retrieved passages $p_{1},\ldots,p_{N}]$ based on their relevance to a query $q$. The listwise ranking approach takes both $q$ and a set of passages as input, and generates a reranked sequence of passage IDs (e.g., [4] [2]…). Due to the constraints of LLMs’ input length, existing listwise ranking approaches typically employ a sliding window strategy that iteratively processes passage subsets. Specifically, this strategy applies a window of size $w$ that moves from the end to the beginning of the passage list with a step size $s$, dynamically promoting relevant passages to higher positions. Following previous studies Sun et al. (2023), we set the total number of passages $N$ to 100, with window size $w$ and step size $s$ set to 20 and 10, respectively.
2.2 Inference Process of CoRanking
Given the passages to rerank, our CoRanking (see Figure 2) first applies the small listwise reranker to pre-rank all the passages using a sliding window strategy. Then, we use the passage order adjuster to rerank the top-20 passages in a listwise manner, so that their order meets the preference of the large listwise reranker. Finally, the large listwise reranker is applied to further rerank the top-20 passages to obtain the final ranking list. This avoids the large listwise reranker’s inefficient sliding window operations on all 100 passages, significantly improving ranking efficiency. The specific prompt template for a large listwise reranker is provided in Appendix A, which is consistent with previous studies Pradeep et al. (2023b); Liu et al. (2024c). In the following, we will illustrate the training details of our small listwise reranker and passage order adjuster.
2.3 Small Listwise Reranker (SLR)
@@
To mitigate the latency and computational overhead of the large listwise reranker (denoted as LLR), we propose using a small listwise reranker (denoted as SLR) to pre-rank the entire passage list. The SLR aims to elevate relevant passages into the final sliding window (i.e., the top-20 passages), thereby reducing the number of LLR’s sliding window iterations. Here we adopt listwise training because we find that it performs better than other approaches (e.g., pointwise training Nogueira and Cho (2019)), and the SFT-trained SLR naturally serves as the initialization for our listwise passage order adjuster, which will be detailed in Section 2.4.
@@
Training Label Conventional listwise reranker training typically employs knowledge distillation Pradeep et al. (2023a, b), which uses a strong teacher model (e.g., GPT-4) to rerank BM25-retrieved top-20 passages, with the teacher-generated rankings serving as supervised fine-tuning (SFT) labels. However, the teacher model can not perform consistently well across all queries, such as long-tail queries, which can result in suboptimal performance. To further improve the SFT label, we propose a human-label-enhanced ranking construction (HRC) method that improves teacher-generated rankings with human labels.
@@
Specifically, we first iteratively rerank BM25-retrieved top-20 passages $M$ times using a powerful teacher model in a listwise manner, where the $(i+1)$-th reranking input is the output ranking from the $i$-th iteration. In this paper, we apply LLR as our teacher model and set the maximum iteration number to 5. This progressive reranking approach is motivated by prior studies Pradeep et al. (2023b, a) which reveal that iteratively reranking enables gradual improvement of ranking lists. Then, we choose the ranking where human-annotated relevant passages are ranked at the top (i.e., NDCG=1) during the iteration process as the gold ranking.
@@
Training Loss After obtaining the gold ranking, we concatenate the query $q$ and the candidate passages $p_{1},\ldots,p_{N}]$ together as the training input $x$ and optimize our SLR by minimizing the standard language modeling loss $\mathcal{L}$:
@@
$x$ $=q\circ\text{[1]}\circ p_{1}\circ\text{[2]}\circ p_{2}\ldots\text{[N]}\circ p_{N},$ (2)
@@
2.4 Passage Order Adjuster (POA)
While our SLR is trained to rank relevant passages at the top of the list, the order of top-ranked passages (i.e., the top-20 passages) may not align well with the order preference of LLR. To address this issue, we employ a reinforcement learning (RL) algorithm to train a passage order adjuster POA based on LLR’s feedback to rerank these top-ranked passages. We adopt the Direct Preference Optimization (DPO) Rafailov et al. (2023) algorithm for its optimization stability. The following sections elaborate on the construction of DPO training data and model optimization.
Training Data The training data of DPO consists of a series of prediction pairs. Each prediction pair contains a winner and a loser prediction, namely, a passage ID list. The reward $R$ for each prediction $L$ is computed by feeding $L$ to LLR and evaluating the ranking metric $M$ of LLR’s output ranking $L^{\prime}$ based on human-annotated relevance label $y$:
@@ 2.4 Passage Order Adjuster (POA)
We apply our trained SLR to rerank BM25-retrieved top-100 passages with a sliding window strategy for training queries and use the top-20 reranked passages as training candidate passages, which aims to align with the input passage distribution encountered by POA during inference. To generate the preference pairs for training POA, we first sample $m$ predictions from SLR by taking the candidates as the input, then evaluate these predictions through Equation 3. After that, the highest-reward list is selected as the winner prediction, while others serve as loser predictions. Next, we pair the winner prediction with all the loser predictions to construct initial training preference pairs.
@@
Optimization Given constructed preference pairs, we optimize POA using the following DPO objective:
@@
where $x$ represents the training input whose format is the same as Equation (2). $y_{w}$, $y_{l}$ represent the winner prediction and loser prediction, respectively. $\pi_{\theta}$ is the policy model to be optimized, and $\pi_{f}$ is the original policy model, which is initialized with our trained SLR. The frozen $\pi_{f}$ acts as a regularizer to maintain generation stability during optimization, and $\beta$ balances preference alignment and regularization towards $\pi_{f}$.
@@ 3.1 Setting
We evaluate our framework on three established information retrieval benchmarks: TREC DL Craswell et al. (2020b), BEIR Thakur et al. (2021), and BRIGHT Su et al. (2024). For the TREC DL benchmark, we adopt the TREC Deep Learning 2019 (DL19) Craswell et al. (2020b) and 2020 (DL20) Craswell et al. (2020a) datasets derived from MS MARCO v1. The BEIR benchmark Thakur et al. (2021) is selected to evaluate cross-domain generalization capabilities. We choose three representative datasets: TREC-Covid, Robust04, and Trec-News. BRIGHT Su et al. (2024) is a reasoning-intensive IR benchmark whose relevance assessment transcends simplistic keyword matching or semantic similarity, instead demanding intentional and deliberate reasoning. We choose Economics, Earth Science, and Robotics datasets for evaluation. The detailed descriptions of each dataset will be provided in Appendix B.
Implementation Details
As for the supervised fine-tuning of our SLR, we use Qwen2.5-3B-Instruct as the backbone model. The training queries consist of a total of 2k queries sampled from the MS MARCO training set. We utilize Qwen2.5-72B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-GPTQ-Int4 as our teacher model to SFT our SLR and also employ it as the LLR model throughout this study unless otherwise specified. As for the RL training of POA, we use our trained SLR for model initialization. We set the threshold $\mu$ as 0.4 and generate about 12k preference pairs from 2k queries for the DPO algorithm. Please see the Appendix D for more implementation details.
Baselines
As for single reranker, we compare our trained SLR with a series of state-of-the-art supervised rerankers, including monoBERT (340M) Nogueira and Cho (2019), monoT5 (3B) Nogueira et al. (2020), RankT5 (3B) Zhuang et al. (2023b), RankVicuna (7B) Pradeep et al. (2023a), RankZephyr (7B) Pradeep et al. (2023b) and $\textbf{RankMistral}_{\text{100}}$ Liu et al. (2024c). The first three are pointwise rerankers trained from human labels in the MS MARCO dataset Nguyen et al. (2016). The last three are listwise rerankers which are distilled from GPT-3.5, GPT-4, and GPT-4o, respectively. To ensure more fair comparison with SLR in model architecture and training data, we introduce baseline Qwen-pointwise, which is a pointwise reranker based on Qwen2.5-3B-Instruct. This model concatenates a query-passage pair as input and outputs a relevance score. We utilize the same gold ranking label as SLR and optimize the model with RankNet loss, which is the same as what  Sun et al. (2023) did. We also report the ranking performance of our SLR and LLR, which rerank all the passages using the sliding window strategy.
As for collaborative rerankers, we introduce three naive collaborative baselines, including monoT5 (3B) + LLR, RankT5 (3B) + LLR, and Qwen-pointwise (3B) + LLR. The three baselines first use a corresponding small reranker (3B) to rerank all the passages and then apply LLR to rerank the top-20 passages.
@@
Table 1: Results (NDCG@10) on TREC, BEIR, and BRIGHT benchmarks. The “Pointwise”, “Sliding”, ’Full’ and “Collaborative” in “Strategy” column represent pointwise ranking, sliding window strategy, full ranking Liu et al. (2024c) and collaborative ranking, respectively. The best and second-best results for the “Single Reranker” part and the “Collaborative Rerankers” part are indicated in bold and underlined, respectively. “Avg.” represents the averaged result of all 8 datasets.